OpenStudy (diyadiya):

If \(\alpha , \beta \) are the roots of \(ax^2-2bx+c=0\) then\(\alpha^3\beta^3+\alpha^2\beta^3+\alpha^3\beta^2\) is :

OpenStudy (diyadiya):

oh ok i got it

OpenStudy (anonymous):

\[ \alpha^3\beta^3+\alpha^2\beta^3+\alpha^3\beta^2 = (\alpha\beta)^3 + (\alpha\beta)^2 ( \alpha + \beta)\]